A $29,000 loan at 12% compounded quarterly is to be repaid by equal quarterly payments over a seven-year term.

a.

What will be the principal component of the sixth payment? (Round your answer to 2 decimal places.)

Principal component of the sixth payment $ .
b.

What will be the interest portion of the twenty-second payment? (Round your answer to 2 decimal places.)

Interest portion $ .
c.

How much will the loan’s balance be reduced by Payments 10 to 15 inclusive? (Round your answer to 2 decimal places.)

Principal paid by Payment 10 to Payment 15 inclusive $ .
d.

How much interest will be paid in the second year? (Round your answer to 2 decimal places.)

Interest paid in the second year $ .
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

SOLVED WITH BA II PLUS CALCULATOR

C W ENG 06:43 01-08-2020 70 35 Х F1898 . B A с ח E D I G J K set C/Y = P/Y = 4 BA II PLUS CALCULATOR 28 [7 X 4] 12.00 N = I/Y

C W ENG 06:43 01-08-2020 70 35 Х G1940 . A B с D E F. G Н J K LOAN BALANCE AFTER 21 PAYMENTS N = 7 [28-21] 1/Y = 12.00 PMT =

C W ENG 06:43 01-08-2020 70 35 Х G1970 . A B с D E F. G Н J K 1947 1948 LOAN BALANCE AFTER 4 PAYMENTS N = 24 [28-4] 1/4 = 12.